how much water must be added to 14 liters of a 90% acid solution to abtain a 40% acid solution?
Answer by stanbon(75887)

how much water must be added to 14 liters of a 90% acid solution to obtain a 40% acid solution?
------------------------------
Equation:
acid + acid = acid
0.90*14 + 0*x = 0.40(14+x)
---
Multiply thru by 100 to get:
90*14 + 0 = 40*14 + 40x
50*14 = 40x
x = (5/4)14
----
x = 17.5 liters (amt. of water that must be added)
===============
Cheers,
Stan H.